OpenStudy (anonymous):

Find the exact value of sin 240°

OpenStudy (zzr0ck3r):

2/3(2PI) = 4PI/3 look on the unit circle

OpenStudy (cwrw238):

this is in 3rd quadrant and = - sin (240-180) = -sin 60

OpenStudy (anonymous):

Well these are my options: negative square root of 2 divided by 2 negative square root of 3 divided by 2 one-half square root of 2 divided by 2

OpenStudy (cwrw238):